OpenStudy (nautika27):

One of Madison's hobbies is to locate hidden geocaches. To find yesterday's cache, Madison was instructed to start at a water fountain at a local park. Then, the directions had her walk 30 yards due west and 16 yards due south. If this was the location of the cache, how far was it from the water fountain where she started? 14 yards23 yards34 yards 46 yards

OpenStudy (nautika27):

@rsmith6559

OpenStudy (nautika27):

@jhonyy9 @Safa102 @Keigh2015 @inkyvoyd @UsernameNotTaken @nikeboi101 @Asad786 @dan815 @rishavraj

OpenStudy (nautika27):

@rsmith6559

OpenStudy (nautika27):

@rsmith6559

OpenStudy (usernamenottaken):

You need to use the Pythagorean formula.

OpenStudy (usernamenottaken):

\[30^2 + 16^2= c^2\]

OpenStudy (usernamenottaken):

\[1156=c^2\]

OpenStudy (usernamenottaken):

c= √1156

OpenStudy (usernamenottaken):

c= 34

OpenStudy (usernamenottaken):

TA DAA